OpenStudy (anonymous):

integral of dx/(Sqrt(-x^2+8x-12))

OpenStudy (anonymous):

-2x^3/3+8x^2/2-12x

OpenStudy (anonymous):

which can be simplified: -2x^3/3 + 4x^2-12x

OpenStudy (anonymous):

sqrt goes to x^(1/2)

OpenStudy (anonymous):

bit lost on the sqrt its been a while. can someone else help?

OpenStudy (anonymous):

1/x goes to ln x

OpenStudy (anonymous):

\[\int\frac{dx}{\sqrt{-x^2+8x-12}}\] Complete the square in the denominator: \[-x^2+8x-12\\ -\left(x^2-8x\right)-12\\ -\left(x^2-8x+16-16\right)-12\\ -\left((x-4)^2-16\right)-12\\ -(x-4)^2+16-12\\ 4-(x-4)^2\] So now you have \[\int\frac{dx}{\sqrt{4-(x-4)^2}}\] Now make a trigonometric substitution. Let \[x-4=2\sin u\\ dx=2\cos u~du\] \[\int\frac{2\cos u}{\sqrt{4-(2\cos u)^2}}~du\]

OpenStudy (anonymous):

Sorry, that last integral should be \[\int\frac{2\cos u}{\sqrt{4-(2\color{red}{\sin u})^2}}~du\]

OpenStudy (anonymous):

why make a sub with trig functions?

OpenStudy (anonymous):

What would you do instead? The procedure you suggested doesn't work.